OpenStudy (anonymous):

Find the excluded value of the rational expression. 2x+6/4x-8

OpenStudy (anonymous):

\[\frac{ 2x+6 }{ 4x-8 }\] \[\frac{ 2(x+3) }{ 4(x-2)}\] \[\frac{ (x+3) }{ (x-2)}\]

OpenStudy (anonymous):

\[\frac{ (x+3) }{ 2(x-2)}\]

OpenStudy (anonymous):

excluded value is x=2

OpenStudy (anonymous):

ty everyone